OpenStudy (angel_kitty12):

Can someone pleaseeee help I've been stuck on this literally all day. I don't get these four questions. Can someone please help? The questions are in two pictures below. Algebra 2 by the way.

OpenStudy (angel_kitty12):

First photo

OpenStudy (angel_kitty12):

Second photo. Please help I've been asking all day but nobodies helped me. Thanks.

OpenStudy (kropot72):

A. The raw score that is 1.5 standard deviations above the mean is found as follows: Required raw score = 56 + 1.5 * 18 = 83 B. Reference to a standard normal distribution table shows that for a z-score of 1.5, the cumulative probability of a raw score being 83 or less is 0.9332. Therefore the probability of a randomly selected student receiving a raw score greater than 83 is given by: P(>83) = 1 - 0.9332 = 0.0668. Can you now calculate how many of the students receive a scaled score greater than 90%? Note: The scaling percentage does not enter the calculation.

OpenStudy (angel_kitty12):

How is 56+1.5*18 83?

OpenStudy (angel_kitty12):

Wait never mind I calculated wrong

OpenStudy (angel_kitty12):

But how did you get 0.9332 from 83?

OpenStudy (angel_kitty12):

@kropot72

OpenStudy (kropot72):

The raw score of 83 was calculated to be 1.5 standard deviations above the mean. Therefore the z-score for 83 must be 1.5. You need to use a standard normal distribution table to find that the cumulative probability for a z-score of 1.5 is 0.9332.